OpenStudy (anonymous):

kara read 30 pages of a book in37 minutes. how long will it take her to read the remaining 247 pages in the book, in minutes? round your answer to the nearest tenth of a minute.

OpenStudy (anonymous):

for this question we set up a ratio. 30/37=247/x

OpenStudy (anonymous):



OpenStudy (anonymous):

thx

OpenStudy (anonymous):

for the help

OpenStudy (anonymous):

Now we write the equation: 30x=37*247 Then divide each side by 30 to solve for x

OpenStudy (anonymous):

np:)

OpenStudy (anonymous):

wait could you tell me the answer cause i have no paper with me

OpenStudy (anonymous):

oh sure! 30.5minutes

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

you were wrong it was 304.6